OpenStudy (anonymous):

Integration by parts: I'm trying to identify u and dv of the following problem: 6x sinx cosx dx. Can someone show me step by step how I do this?

OpenStudy (dumbcow):

i would first simplify it by using the double angle identity \[\sin(2x) = 2 \sin x \cos x\] \[\rightarrow 3\int\limits x \sin(2x) dx\] u = x .......... dv = sin(2x) du = dx ........ v = -cos(2x)/2

OpenStudy (anonymous):

Ok, thank you.

OpenStudy (dumbcow):

yw